OpenStudy (anonymous):

Write an equation for the line that contains the point (-2,9) and is parallel to the line y=2x+6.

OpenStudy (anonymous):

A. Y=-(1/2)x+8 B. Y=2x-7 C. Y= 2x+11 D. Y=2x+13